Prev Next

SW&IT Meeting Minutes 2020-05-11

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20200504
LR20200504 - Bec, Matthieu 40.0
LR20200504 - Holmberg, Carl 40.0
LR20200504 - Edmeades, David 0.0
LR20200504 - Sola, Igor 40.0
LR20200504 - Hooper, Stephen 80.0
LR20200504 - Zhang, Yang 0.0
LR20200504 - Zhang, Xianyu 40.0

News from management

  • REWG - Re-Entry working group. Pat, Mike, David, Laura, Christian

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • printer for J.Wiese
  • CR restart May 18 - include archive VM/NAS
  • FOSWiki landing page

Team reports

PK

https://github.com/mermaid-js/mermaid and https://stackedit.io rocks for editing GitHub MD with graphs.

MB

  • Last week
    • OC, Management
    • restart 3-options planning
    • DMS
      • MS Edge / Firefox: grid-layout and <img/> animations, many fixes
    • AO PB/SOUL/ARGOS/LN meetings, draft scenarii presentations
    • AO weekly sessions
    • Safety trainings complete
  • This week
    • remote operation requirement document (due this week)
    • DMS
      • weather loss, websockets, AdSec, etc.
      • LBC stiched png preview
    • SHARK-VIS new IIF document and clarify DD entries for SOUL
    • follow up student ad
    • cobbler / kickstart, ims f30

DE

  • Veeam quote followup
  • News from Zoom: will force update of client to v. 5 by end of month
  • oVirt tasks
    • Added VMWare as provider
    • Test VM transfer
    • Mixed results
      • Machine boots
      • LVM very unhappy
  • KVM troubleshooting
    • Found some options; more detail with IT
    • "Java-free" for certain values of "free"
  • Cisco training purchase workaround
  • Passive half of Synology in Tucson is unhappy; suggest reboot
  • Jim printer
    • As in email, I have reservations about how likely it is that Jim will be able to get a new printer working. We could remote in for some setup, but there are plenty of issues that need hands on to solve.
    • Cheap option is inkjet, with all of the hassles and long term expenses that that brings
    • Laser is quite expensive, up to Bucki to decide.
  • Rofin/laser cutter computer order ongoing; waiting for reply from Robert on how he'd like to proceed.
    • At least we didn't need Peter to go to the machine shop to grab the configuration data
  • Purchase lucix upgrade hardware so we have it on hand?

YZ

  • Previous:
    • AO training daily
    • Some debugging with AO-Cravat Strehl code.
    • modsAlign refactoring.
    • TMS information review.
    • MODS WCS header discussion with Olga.
  • This week:
    • Continue AO-Cravat work (some Strehl code debugging and hopefully discussion with Al and Steve. Start to think about document generation).
    • modsAlign refactoring.
    • TMS information review and discussion.
    • Start to work on a Python 3 infrastructure for MODS header editing.
    • Review Olga's current info related to WCS calculation.
    • Check with Robert on TCS/IIF doc VCAN release?

XZ

  • Last week
    • Merge UAO to SOUL
    • Setup and test the different AO versions on the VM simulator
    • AO PB/SOUL/ARGOS/LN discussion
    • PB SW/HW discussion
    • Review SOUL slopeBCU document
    • Meetings
  • This week
    • troubleshooting of SX ASM issue
    • working on PB action items
    • Merge UAO to SOUL
    • PB SW/HW

SH

  • abell moved to ovirt, John says it is working
  • web moved to ovirt
  • wiki moved to ovirt, upgraded to Fedora 31
  • finished wiring in Tucson CR
  • new private network -- need to discuss this with UITS
  • web camera access for MGIO
  • three more monitors for ROWG
  • SSL for LBTC

next:

  • ROWG monitor
  • DNS on mountain and in Tucson
  • termserv (what do we do)
    • vault has the same kind of issue
  • move at least three developer machines from vm2.tucson

CH

Last Week:
  • DMS: changes imonnit description generator and ingest script.
  • Safety training
This Week:
  • AGW: support Juan Pablo and John Morris for testing.
  • DMS: complete imonnit ingest changes.
  • OA Control: Research software changes needed for encoder support binary output.
  • Safety training

IS

Last week:
  • LUCI
    • AO Training
    • Build LUCI environment in a virtualbox vm Centos7.
    • Still some work to do particularly in the database. This work is useful because:
      • it will allow development in LUCI locally avoiding the VPN.
      • We might need to reinstall LUCI after David's upgrade in the LUCI machine.
      • Help me in case I have get the lucix backup environment going.
This week:
  • LUCI
    • finish the VM work.
    • work on #IT7952 an RTD/Aladin GUI issue.
    • Look into LUCI WebIO device configuration.
  • PIT
    • If I find some time, work on the PIT admin tool backend.
  • OT
    • work with Andrew, if available, on current test OT release issues
Topic revision: r15 - 17 May 2020, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback